Score: 0 of 1 pt output has come from other c s. In 1998, the total demand for wheat was Q 3,244-283P Qp 1700-107P Q2 1,944 + 207P U.S are concened about this drop in export demand What happens to the free-market price of wheat in the United States? The free-market price of wheat in the United States after the drop in export demand is $ (Enter your response rounded to two decimal places)
Answer #1

Given total demand, Q = 3244 - 283P, and domestic demand, Qd = 1700 - 107P,

subtract and find export demand,

Qe = 1544 - 176P.

The initial market equilibrium price is found by setting total demand equal to supply:

3244 - 283P = 1944 + 207P,

or P = $2.65.

The best way to handle the 20 percent drop in export demand is

The new export demand is

(1- 0.2)Qe= 0.8(1544-176P)= 1235.2 - 140.8P

Total demand becomes

QD = Qd + 0.8Qe

= 1700 - 107P + 1235.2 - 140.8P

= 2935.2 -247.8P

Equating total supply and total demand,

1944 + 207P = 2935.2 - 247.8P

= 454.8 P = 991.2

P = $2.18